OpenStudy (anonymous):

Evaluate 9 − 3 • 2 + 4 a.) 6 b.) 7 c.) 15 d.) 16

OpenStudy (anonymous):

D. 16

OpenStudy (anonymous):

9-3= 6 6 * 2 = 12 12 + 4 = 16

OpenStudy (noseboy908):

PEMDAS. Parenthesis, Exponents, Multiplication/Division, and lastly Addition/Subtraction.